betty is thinking of two consecutive integer whose sum is 41. Let x represnt the smaller unknown integer.
a. How could you represnt the larger unknown integer in terms of x?

Answer by bucky(2189)

Since the integers are consecutive, if the smaller integer is x then the next integer
has to be one greater than x. So we can represent the two integers as x and x+1. Since
the sum of the two integers is 41 we can write:
.
x + (x + 1) = 41
.
Remove the parentheses on the left side to make the equation:
.
x + x + 1 = 41
.
Combine the two x's to get:
.2x + 1 = 41
.
Subtract 1 from both sides to eliminate the 1 on the left side. This results in:
.
2x = 40
.
Divide both sides by 2 to solve for x and the result is:
.
x = 40/2 = 20
.
Since the two integers are x and x + 1, the integers are 20 and 20 + 1 or 20 and 21.
.
If you add 20 and 21, you find that the sum of these two consecutive integers equals 41,
just as the problem required.
.
Hope this helps you to understand the problem a little better ...